OpenStudy (anonymous):

differentiate G(t)=t^3cost

OpenStudy (amistre64):

I am assuming that G(t) = (t^3)(cos(t)); use product rule to derive. Dx(f(x)) = L'R + LR' Dx(G(x)) = (3t^2)(cos(t)) + (t^3)(-sin(t)) G'(x) = (3t^2)(cos(t)) - (t^3)(sin(t)) G'(x) = (t^2) [3cos(t) - t sin(t)] thats the best I can do with it :)

OpenStudy (amistre64):

just for correction its: G'(t) = .....

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can u answer any chemistry questions

OpenStudy (anonymous):

?

OpenStudy (amistre64):

I dont know much about chemistry, sorry...

OpenStudy (anonymous):

np